Use the Theorem of Pythagoras a2 +b2 = c2 to calculate the following: (7)
A point on the ground is 6 m from a pole that is 12 m tall. Calculate the length of a straight line between this point and the top of the pole.

Answer by nerdybill(5405)

A point on the ground is 6 m from a pole that is 12 m tall. Calculate the length of a straight line between this point and the top of the pole.
.
Let x = length of a straight line between this point and the top of the pole
then applying Pythagorean Theorem we get:
.
6^2 + 12^2 = x^2
36 + 144 = x^2
180 = x^2
sqrt(180) = x
13.416 m = x